# Gravitational waves from strong first order phase transitions

### v1

@@ I Introduction
The detection of gravitational waves from black hole and neutron star mergers [1, 2, 3, 4] has inaugurated the age of gravitational wave astronomy. In addition, the detection of signals compatible with a nanohertz stochastic gravitational wave background (SGWB) by the NANOGrav [5] and EPTA [6] collaborations, has renewed the focus on gravitational waves as a window through which to study early universe phenomena [5].
A possible source of primordial gravitational waves would come from anisotropic stresses generated by first-order phase transitions. Although the Standard Model of particle physics has only cross-overs at the QCD [7] and electroweak [8, 9] transitions, and would therefore generate no visible imprint on the SGWB, many Beyond the Standard Model theories allow first-order phase transitions at - and above - the electroweak scale (see [10] for an overview, or for some examples [11, 12, 13, 14, 15, 16]) or at the QCD scale [17, 18]; possibly also in a dark sector (see e.g. [19, 20, 21, 22]).
@@
First-order phase transitions at energy scales between $100$ GeV and $1$ TeV are expected to generate SGWB imprints in the mHz frequency range, making them detectable by the future Laser Interferometer Space Antenna (LISA) [25, 10, 26]. Pulsar Timing Array experiments could detect gravitational waves from phase transitions at lower energy scales of 100100100\,100MeV – 111\,1GeV [27] (see also the latest Pulsar Timing Array constraints in the Erratum of Ref. [28]).
@@
Characterising non-linear hydrodynamic processes is difficult because of the computational requirements of 3D simulations. One needs to accurately resolve multiple length scales all the way from the bubble wall thickness to the size of sound horizon, which requires a large and finely discretised lattice. Additionally, the whole simulated system must evolve for a sufficiently long time to allow the non-linear processes to operate. In the case of transitions proceeding by detonations, where there is no interaction of the pressure waves surrounding the expanding bubbles before they collide, significant savings in memory and time requirements can be found by modelling the order parameter dynamics in the advancing phase front as a discontinuous change in the equation of state, in so-called Higgsless simulations [33, 35]. However, in deflagrations, where the pressure waves in front of expanding bubbles interact and slow down the bubble walls [32], Higgless simulations miss important physical effects. Still, the latest results from these simulations (Ref. [35]) highlight the importance of understanding kinetic energy decay for calculations of the gravitational wave spectrum, especially for strong first order phase transitions.
@@
The UETCs also show the effect of non-linearities. We see the expected Gaussian decorrelation with time difference in the vortical flow of the deflagration, as predicted by the Kraichnan random sweeping model [48], in which short-distance modes are “swept” by the average velocity at the dominant scale of the flow. We also see the sinusoidal behaviour of the compressional mode in both the deflagration and detonation, expected from acoustic oscillations. However, the sinusoid is modulated by a Gaussian envelope, and the correlations die out after a few oscillations. Furthermore, the oscillation period for a given wavenumber is faster than would be predicted for waves travelling at the sound speed, indicating that the decorrelation is caused by the motion of the shocks. The time decorrelation of the subdominant vortical mode of the deflagration is also interesting, showing a combination of Gaussian and oscillatory behaviour. We can model this in terms of random sweeping by acoustic oscillations.
@@
Our simulations are long enough to see the square of the kinetic energy decline by $80\%$, and hence the rate of gravitational wave production, which is proportional to the square of the kinetic energy, is then significantly lower by the end. At frequencies higher than the peak, the gravitational wave power spectrum saturates, and we are able to make a firm prediction of the final power spectrum. The gravitational wave power is still growing around the peak frequency, but we can extrapolate the observed growth of the power spectrum to make a more uncertain, but still useful prediction. We find that the total gravitational power can be expressed in terms of a dimensionless efficiency factor of O($10^{-2}$), indicating that a good estimate of the effective lifetime of the flow is given by taking it to last for one shock decay time, computed from the mean bubble spacing divided by the initial RMS velocity. A similar efficiency factor has also been observed in smaller Higgsless simulations [35].
@@
This paper is divided as follows: a summary of the acoustic generation of gravitational waves and key observables (Section II); methodology of the simulation setup, with additional details added in appropriate appendices (Section III); the main results (Section IV); analysis of the results in terms of a model of acoustic production (Section V); and a conclusion summarising the manuscript and outlining possible next steps (Section VI).
@@ II First-order phase transitions
We begin by reviewing the thermodynamics and hydrodynamics of a thermal first-order phase transition as might have taken place in the early universe (see e.g. [50, 23]). The transition is driven by a scalar field, for example the Higgs field, or an effective scalar field, functioning as an order parameter. This is coupled to other fields, such as those of the Standard Model. The excitations of these fields can be approximated as a relativistic fluid whose equation of state depends on the order parameter.
@@
As the bubbles expand, excess free energy in the metastable phase is transferred to shells of moving and heated fluid via the coupling between the order parameter and the fluid. This effective dissipative coupling models the friction: this arises from a departure from equilibrium as particles change their masses when crossing the wall [51, 52]. The coupling dictates the expansion speed of the bubble wall. If this coupling is negligible and therefore the frictional force on the fluid is insufficient to stop the walls from becoming ultrarelativistic, the transition is known as a “runaway” [51, 52]. This scenario has been simulated in Refs. [53, 54, 55, 56], but is not considered here.
If the speed is subsonic, the hydrodynamic solution is a compression wave in front of the bubble wall with a leading shock, known as a deflagration [57]. For walls which are supersonic, but whose speed is lower than a critical threshold (known as the Chapman-Jouguet speed), a rarefaction wave develops behind the wall, while the compression wave in front still exists. Past this limiting speed, the wall catches up with the leading shock, and thus for faster wall speeds the fluid disturbance is a detonation, a pure rarefaction wave behind the bubble wall.
@@ II.1 Effective theory of a first order phase transition
where $\phi_{b}$ is the order parameter at the $\phi\neq 0$ minimum of $V_{0}$, and $\Delta V_{0}=V_{0}(0)-V_{0}(\phi_{b})$. Hence $V(0,T_{\text{c}})=V(\phi_{b},T_{\text{c}})$, as required. The consequences of this thermodynamic model for phase transition simulations was briefly examined in Ref. [60].
@@
where $T_{\text{n}}$ is the nucleation temperature, defined as the temperature at the maximum volume-averaged bubble nucleation rate. In exponential nucleation, where the nucleation rate density increases exponentially with time after the critical temperature is reached, this coincides with there being a volume fraction of $1/e$ in the broken phase. In the simulations described herein, all bubbles are nucleated simultaneously at $t=0$.
@@
$H_{\text{n}}^{2}=\frac{8\pi G}{3}\bar{\epsilon}_{0}\;.$ (8)
@@
Now that the thermodynamics of the phase transition have been described, we can move towards the equations of motion of field-fluid system. We use the metric convention ($+$)(-\,+\,+\,+)( - + + + ). If the fluid and the field are uncoupled, we can define corresponding energy-momentum tensors
@@
Hence Parseval’s theorem is expressed as $\int\mathrm{d}k\,\mathcal{P}_{Y}(k,t)/k=\langle Y^{2}({\mathbf{x}},t)\rangle$. We we also refer to the spectral density as an equal time correlator (ETC). Note that the power spectrum has the same dimensions as $Y^{2}$, that is $[{\mathcal{P}}_{Y}]=[Y]^{2}$, while $[P_{Y}]={\sf L}^{3}[Y]^{2}$.
@@
where $\Lambda_{ijkl}$ is the transverse-traceless projector,
@@ II.3 Analytical expectations
The vortical modes should develop turbulence and evolve towards a Kolmogorov spectrum, which for the power spectrum as we define it is $k^{-2/3}$ [31]. However, this takes several eddy turn-over times. As far as we are aware, there are no theoretical predictions for the initial power spectrum of the vortical modes laid down by bubble collisions.
@@ III Simulation setup
$\begin{split}-\ddot{\phi}+\nabla^{2}\phi-\frac{\partial V}{\partial\phi}=\eta\gamma(\dot{\phi}+v^{i}\partial_{i}\phi)\;,\end{split}$ (30)
and for the fluid we obtain,
$\begin{split}\dot{E}+\partial_{i}(Ev^{i})&+p[\dot{\gamma}+\partial_{i}(\gamma v^{i})]-\frac{\partial V}{\partial\phi}\partial_{i}\phi\\ &=\eta\gamma^{2}(\dot{\phi}+v^{i}\partial_{i}\phi)^{2}\;,\end{split}$ (31)
$\begin{split}\dot{Z}+\partial_{j}(Z_{i}v^{j})&+\partial_{i}p+\frac{\partial V}{\partial\phi}\partial_{i}\phi\\ &=-\eta\gamma(\dot{\phi}+v^{j}\partial_{j}\phi)\partial_{i}\phi\;,\end{split}$ (32)
@@
Our simulation grid has spacing $\delta x=1.0T_{c}^{-1}$, with timesteps of size $\delta t=0.2T_{c}^{-1}$, a choice which in past works [32] demonstrated reasonable energy conservation (see Appendix B), captured the dynamical features of bubble collision and was sufficient to accurately describe the gravitational wave power spectra.
The phase transition is initiated by creating 64 Gaussian bumps in $\phi$ whose centre is in the broken phase $\phi_{b}=2.0T_{c}$ at random positions on the simulation grid, while avoiding overlaps. This results in a mean bubble spacing of $R_{*}=1024$. These bumps are large enough that they are guaranteed to expand, and during their evolution they relax towards the a self-similar velocity and enthalpy profile (see Appendix C). The fluid, initially at rest, is driven by the pressure of expanding scalar field bubbles to form outward-moving fluid shells.
@@
(e) |×$|\nabla\times{\mathbf{v}}|(x,y,z=0)$
(f) |×$|\nabla\times{\mathbf{v}}|(x,y,z=0)$
Figure 1: Slices through simulations. Left: detonation at time $t=5000/T_{c}$. Right: deflagration at time $t=7700/T_{c}$. Top panels show the magnitude of the 3-velocity $|{\mathbf{v}}|$, middle panels the magnitude of the divergence of spatial components of the 3-velocity $|\nabla\cdot{\mathbf{v}}|$, and the bottom panels showcase the magnitude of the curl |×$|\nabla\times{\mathbf{v}}|$.
We now list the main observables, which consist of volume-average quantities, power spectra, and unequal-time correlators.
We have two estimators of the average wall speed [60]. The first is obtained from the volume average kinetic and gradient energies of the scalar field,
@@
$v_{\text{w},\phi}=\bigg{(}\frac{e_{\dot{\phi}}}{e_{D}}\bigg{)}^{1/2}$ (35)
The second comes from the rate of change of the total number of points in the broken phase, along with the estimated surface area of walls,
@@
computed via total number of sites where the scalar field changes sign (ie. sites where $\phi_{i}\phi_{i+1}<0$).
@@
Simulations start at time $t=0$ and run to final times $t_{\text{end}}=4990/T_{\text{c}}$ (detonation) and $8000/T_{\text{c}}$ for the main deflagration run. A separate run with $t_{\text{end}}=5880/T_{\text{c}}$ was used to measure UETCs in the deflagration.
@@
Phase transition Interval [$T_{\text{c}}^{-1}$] $t_{\text{ref}}T_{\text{c}}$
@@
A selection of slices from each of the simulations is shown in Fig. 1. They show pseudocolour plots of magnitude of 3-velocity $|{\mathbf{v}}|$, divergence of 3-velocity $|\nabla\cdot{\mathbf{v}}|$ and its curl |×$|\nabla\times{\mathbf{v}}|$. The lower two are presented in logarithmic scale to highlight the presence of relevant non-linearities (shocks, which appear as bright streaks, and vorticity in the form of eddies). These are presented for times $t=5000/T_{c}$ in the detonation, and $t=7700/T_{c}$ for the deflagration. An animation for the deflagration can be found at the following link [69].
A noteworthy feature in the detonation is the appearance of three-way junctions of shocks or $\lambda$-configurations [70]. These appear between shocks colliding at less than a certain critical angle, which is a function of the Mach number and the adiabatic index [70, 71], and can be seen in the laboratory and in astrophysical flows [72]. The $\lambda$-configurations already indicate non-linear interactions in the fluid flow.

### v2

@@ I Introduction
The detection of gravitational waves from black hole and neutron star mergers [4, 1, 2, 3] has inaugurated the age of gravitational wave astronomy. In addition, the detection of signals compatible with a nanohertz stochastic gravitational wave background (SGWB) by the NANOGrav [7] and EPTA [10] pulsar timing array collaborations, has renewed the focus on gravitational waves as a window through which to study early universe phenomena [7].
A possible source of primordial gravitational waves would come from anisotropic stresses generated by first-order phase transitions. Although the Standard Model of particle physics has only cross-overs at the QCD [19] and electroweak [61, 62] transitions, and would therefore generate no visible imprint on the SGWB, many Beyond the Standard Model theories allow first-order phase transitions at - and above - the electroweak scale (see Ref. [26] for an overview, or for some examples [76, 27, 16, 42, 15, 73]) or at the QCD scale [11, 82]; possibly also in a dark sector (see e.g. [79, 5, 23, 34]).
@@
First-order phase transitions at energy scales between $100$ GeV and $1$ TeV are expected to generate SGWB imprints in the mHz frequency range, making them detectable by the future Laser Interferometer Space Antenna (LISA) [9, 26, 14]. Pulsar timing array experiments could detect gravitational waves from phase transitions at lower energy scales of 100100\,MeV – 11\,GeV [84] (see also the latest pulsar timing Array constraints in the Erratum of Ref. [8]).
@@
Characterising non-linear hydrodynamic processes is difficult because of the computational requirements of 3D simulations. One needs to accurately resolve multiple length scales all the way from the bubble wall thickness to the size of sound horizon, which requires a large and finely discretised lattice. Additionally, the whole simulated system must evolve for a sufficiently long time to allow the non-linear processes to operate. In the case of transitions proceeding by detonations, where there is no interaction of the pressure waves surrounding the expanding bubbles before they collide, significant savings in memory and time requirements can be found by modelling the order parameter dynamics in the advancing phase front as a discontinuous change in the equation of state, in so-called Higgsless simulations [58, 25]. However, in deflagrations, where the pressure waves in front of expanding bubbles interact with each other [37], Higgless simulations miss important physical effects such as slowing down of the bubble walls. Still, the latest results from these simulations (Ref. [25]) highlight the importance of understanding kinetic energy decay for calculations of the gravitational wave spectrum, especially for strong first order phase transitions.
@@
The UETCs also show the effect of non-linearities. We see the expected Gaussian decorrelation with time difference in the vortical flow of the deflagration, as predicted by the Kraichnan random sweeping model [46]. In this model, short-distance modes are “swept” by the average velocity at the dominant scale of the flow. We also see sinusoidal behaviour of the compressional mode in both the deflagration and detonation, as would be expected from acoustic oscillations. However, the sinusoid is modulated by a Gaussian envelope, and the correlations die out after a few oscillations. Furthermore, the oscillation period for a given wavenumber is faster than would be predicted for waves travelling at the sound speed, indicating that the decorrelation is caused by the motion of the shocks. The time decorrelation of the subdominant vortical mode of the deflagration is also interesting, showing a combination of Gaussian and oscillatory behaviour. We can model this in terms of random sweeping by acoustic oscillations.
@@
Our simulations are long enough to see the square of the kinetic energy decline by $80\%$, and hence the rate of gravitational wave production, which is proportional to the square of the kinetic energy, is then significantly lower by the end. At frequencies higher than the peak, the gravitational wave power spectrum saturates, and we are able to make a firm prediction of the final power spectrum. The gravitational wave power is still growing around the peak frequency, but we can extrapolate the observed growth of the power spectrum to make a more uncertain, but still useful prediction. We find that the total gravitational power can be expressed in terms of a dimensionless efficiency factor of $\mathrm{O}(10^{-2})$, indicating that a good estimate of the effective lifetime of the flow is given by taking it to last for one shock decay time, computed from the mean bubble spacing divided by the initial RMS velocity. A similar efficiency factor has also been observed in smaller Higgsless simulations [25].
@@
This paper is divided as follows: a summary of the acoustic generation of gravitational waves and key observables (Section II); methodology of the simulation setup, with additional details added in appropriate appendices (Section III); the main results (Section IV); analysis of the results in terms of a model of acoustic production (Section V); and a conclusion summarising the manuscript and outlining possible next steps (Section VI). Scripts used for the figures in this manuscript, and the accompanying data (sans simulation slices) are available at [30] and [28], respectively.
@@ II First-order phase transitions
We begin by reviewing the thermodynamics and hydrodynamics of a thermal first-order phase transition as might have taken place in the early universe (see e.g. [72, 48, 12]). The transition is driven by a scalar field, for example the Higgs field, or an effective scalar field, functioning as an order parameter. This is coupled to other fields, such as those of the Standard Model. The excitations of these fields can be approximated as a relativistic fluid whose equation of state depends on the order parameter.
@@
As the bubbles expand, excess free energy in the metastable phase is transferred to shells of moving and heated fluid via the coupling between the order parameter and the fluid. This effective dissipative coupling models the friction as particles change their masses when crossing the wall [18, 17], which causes a departure from equilibrium. The coupling dictates the expansion speed of the bubble wall. If this coupling is negligible and therefore the frictional force on the fluid is insufficient to stop the walls from becoming ultrarelativistic, the transition is known as a “runaway” [18, 17]. This scenario has been simulated in Refs. [36, 66, 35, 68], but is not considered here.
If the wall speed is subsonic, the hydrodynamic solution is a compression wave in front of the bubble wall with a leading shock, known as a deflagration [65]. For walls which are supersonic, but whose speed is lower than a critical threshold (known as the Chapman-Jouguet speed), a rarefaction wave develops behind the wall, while the compression wave in front still exists. Past this limiting speed, the wall catches up with the leading shock, and thus for faster wall speeds the fluid disturbance is a detonation, a pure rarefaction wave behind the bubble wall.
@@ II.1 Effective theory of a first order phase transition
where $\phi_{b}$ is the order parameter at the $\phi\neq 0$ minimum of $V_{0}$, and $\Delta V_{0}=V_{0}(0)-V_{0}(\phi_{b})$. Hence $V(0,T_{\text{c}})=V(\phi_{b},T_{\text{c}})$, as required. By construction then, $\Delta V_{0}=V_{c}$. The consequences of this thermodynamic model for phase transition simulations were briefly examined in Ref. [38].
@@
where $T_{\text{n}}$ is the nucleation temperature, which we define as the temperature at the maximum volume-averaged bubble nucleation rate. [^1] Other choices of notation exist. A popular definition of the nucleation temperature is the temperature at which on average one bubble is nucleated per Hubble volume. Furthermore, some authors denote by $T_{*}$ what we call $T_{\text{n}}$ in the present work. In exponential nucleation, where the nucleation rate density increases exponentially with time after the critical temperature is reached, this coincides with there being a volume fraction of $1/e$ in the broken phase. In the simulations described herein, all bubbles are nucleated simultaneously at $t=0$.
@@
$H_{\text{n}}^{2}=\frac{8\pi G}{3}\bar{\epsilon}_{0},$ (8)
@@
Now that the thermodynamics of the phase transition have been described, we can move towards the equations of motion of the field-fluid system. We use the metric convention (+++)(-\,+\,+\,+). If the fluid and the field are uncoupled, we can define corresponding energy-momentum tensors
@@
Hence Parseval’s theorem is expressed as $\int\mathrm{d}k\,\mathcal{P}_{Y}(k,t)/k=\langle Y^{2}({\mathbf{x}},t)\rangle$. We we also refer to the spectral density as an equal time correlator (ETC). Note that the power spectrum has the same dimensions as $Y^{2}$, that is ${\mathcal{P}}_{Y}]=[Y]^{2}$, while $P_{Y}]={\sf L}^{3}[Y]^{2}$.
@@
where $\Lambda_{ij,kl}$ is the transverse-traceless projector,
@@ II.3 Analytical expectations
The vortical modes should develop turbulence and evolve towards a Kolmogorov spectrum, which for our definition of the velocity power spectrum is $k^{-2/3}$ [13]. However, this takes several eddy turn-over times. As far as we are aware, there are no theoretical predictions for the initial power spectrum of the vortical modes laid down by bubble collisions.
@@ III Simulation setup
$\ddot{\phi}+\nabla^{2}\phi-\frac{\partial V}{\partial\phi}=\eta\gamma(\dot{\phi}+v^{i}\partial_{i}\phi)\;,$ (30)
while for the fluid we obtain,
$\dot{E}+\partial_{i}(Ev^{i})+p[\dot{\gamma}+\partial_{i}(\gamma v^{i})]-\frac{\partial V}{\partial\phi}\partial_{i}\phi\\ =\eta\gamma^{2}(\dot{\phi}+v^{i}\partial_{i}\phi)^{2}$ (31)
and
$\dot{Z}+\partial_{j}(Z_{i}v^{j})+\partial_{i}p+\frac{\partial V}{\partial\phi}\partial_{i}\phi\\ =-\eta\gamma(\dot{\phi}+v^{j}\partial_{j}\phi)\partial_{i}\phi\;,$ (32)
@@
Our simulation grid has spacing $\delta x=1.0T_{c}^{-1}$, with timesteps of size $\delta t=0.2T_{c}^{-1}$. In previous works (e.g. [37]) these choices demonstrated reasonable energy conservation (see Appendix B) while capturing the dynamical features of bubble collision and the variation of gravitational wave energy density. In particular, this choice was seen to be sufficient to accurately describe the gravitational wave power spectra for our chosen values of $v_{w}$ and $\alpha_{n}$.
The phase transition is initiated by creating 64 Gaussian bumps in $\phi$ whose centre is in the broken phase $\phi_{b}=2.0T_{c}$ at random positions on the simulation grid, while avoiding overlaps. This results in a mean bubble spacing of $R_{*}=1024$. These bumps are large enough that they are guaranteed to expand, and during their evolution they relax towards a self-similar velocity and enthalpy profile (see Appendix C). The fluid, initially at rest, is driven by the pressure of expanding scalar field bubbles to form outward-moving fluid shells.
@@
(e) $|\nabla\times{\mathbf{v}}|(x,y,z=0)$
(f) $|\nabla\times{\mathbf{v}}|(x,y,z=0)$
Figure 1: Slices through simulations. Left: detonation ($\alpha_{n}=0.67$, $v_{w}=0.92$) at time $t=5000/T_{c}$. Right: deflagration ($\alpha_{n}=0.50$, $v_{w}=0.44$) at time $t=7700/T_{c}$. Top panels show the magnitude of the 3-velocity $|{\mathbf{v}}|$, middle panels the magnitude of the divergence of spatial components of the 3-velocity $|\nabla\cdot{\mathbf{v}}|$, and the bottom panels showcase the magnitude of the curl $|\nabla\times{\mathbf{v}}|$.
We now list the main observables, which consist of volume-averaged quantities, power spectra, and unequal-time correlators.
We have two estimators of the average wall speed [38]. The first is obtained from the volume-averaged kinetic and gradient energies of the scalar field,
@@
$v_{w,\phi}=\bigg(\frac{e_{\dot{\phi}}}{e_{D}}\bigg)^{1/2}$ (35)
The second comes from the rate of change of the total number of points in the broken phase,
@@
where $A$ is the total number of sites where the scalar field changes sign (ie. sites where $\phi_{i}\phi_{i+1}<0$).
@@
Simulations start at time $t=0$ and run to final times $t_{\text{end}}=4990/T_{\text{c}}$ (detonation) and $8000/T_{\text{c}}$ for the main deflagration run. A separate run with the same initial conditions and $t_{\text{end}}=5880/T_{\text{c}}$ was used to measure UETCs in the deflagration.
@@
Phase transition Interval $T_{\text{c}}^{-1}$] $t_{\text{ref}}T_{\text{c}}$
@@
A selection of slices from each of the simulations is shown in Fig. 1. They show pseudocolour plots of magnitude of 3-velocity $|{\mathbf{v}}|$, divergence of 3-velocity $|\nabla\cdot{\mathbf{v}}|$ and its curl $|\nabla\times{\mathbf{v}}|$. The latter two are presented in logarithmic scale to highlight the presence of relevant non-linearities (shocks, which appear as bright streaks, and vorticity in the form of eddies). These are presented for times $t=5000/T_{c}$ in the detonation, and $t=7700/T_{c}$ for the deflagration. An animation for the deflagration is available at [29].
A noteworthy feature in the detonation is the appearance of three-way junctions of shocks, termed $\lambda$-configurations [31]. These appear between shocks colliding at less than a certain critical angle, which is a function of the Mach number and the adiabatic index [31, 41], and can be seen in the laboratory and in astrophysical flows [47]. The $\lambda$-configurations already indicate non-linear interactions in the fluid flow.